# A Quarter of a Century of Neuromorphic Architectures on FPGAs - an Overview

### v2

@@ 3.8. Class 7 (Traits: fully parallel, collocated computation and memory, asynchronous network update)
Early implementations. Allen et al. (2005) (Allen et al., 2005) presented a platform for pattern recognition in olfactory applications. The system supported IF neuron model with weights uniformly set to 1, so the output spike train of the neuron had a firing rate equal to the average rate of the input spike trains. We deduced that architecture-wise the system was an array of PEs - one for every neuron. The authors reported eight neurons with 1024 input synapses implemented on the Xilinx Virtex-II Pro device. Girau et al. (2006) (Girau and Torres-Huitzil, 2006) presented an architecture for a Local Excitatory Global Inhibitory Oscillator Network (LEGION)(Wang and Terman, 1995), based on IF neurons, for image inference.
Figure 10. Simplified diagram of a Class 7 NMA.
@@
Architecture-wise, due to the inherent nature of the LEGION-type SNNs, the implementation comprised a 2D array of PEs. Neurons were interconnected with excitatory connections, and there was a single global inhibitor connected to all neurons. The size of the array was equal in size to the size of the image under classification. Cassidy et al. (2007) (Cassidy et al., 2007) presented a 2D array of LIF neurons capable of realizing Spatio-Temporal Receptive Fields (STRFs), neural parameter optimization algorithms and STDP. The architecture comprised 32 PEs set up as an array, connected to the weight and delay memory. The datapath was time-multiplexed to allow for parallel processing and serial communication. Every neuron was connected to its four neighbors. The system was implemented on Xilinx Spartan-3 device and was clocked at 50 MHz.
@@
Early modern implementations. Bonabi et al. (2021) (Bonabi et al., 2012) the authors presented a network of 16 HH neurons on a Xilinx Virtex-7 device. The neuronal dynamics were approximated with the CORDIC algorithm and step-by-step time integration. The authors focused on neuronal pools, i.e., groups of interconnected neurons (in A2A fashion) with specific functions, which could be either diffusive or localized. Dean et al. (2014) (Dean et al., 2014) presented DANNA - an NMA that comprised an array of multi-purpose elements that could be configured to perform a function of a neuron, synapse or fan-in/fan-out connection. The architecture was a 2D grid of elements, which be neurons, synapses or fan-out elements. The architecture supported the LIF neuron model with axonal delays. The synapse model supported long-term potentiation and depression for the weights. The network topology was obtained by performing the genetic algorithm for finding the optimal implementation. The authors tested the system [^10] There were many examples of the use cases for this architecture, including NeoN(Mitchell et al., 2017), which was a neuromorphic control for autonomous robotic navigation. with a randomly connected network of 10000 neurons, implemented on a Xilinx Virtex-7 device with a 9b fixed-point number representation, clocked at eight megahertz, while the elements was clocked at 64 MHz. The system was also updated in 2018 and presented as DANNA 2(Mitchell et al., 2018). Nanami et al. (2016) (Nanami and Kohno, 2016) presented a modification of an existing Digital Spiking Silicon Neuron (DSSN)(Kohno and Aihara, 2007) idea. The proposed architecture used 16 DSSNs connected to 16 synapse units, each supporting 256 external synapses - all operating at 10x real-time speed. The number representation was 18b. There were no implementation details provided in the paper, so it is difficult to draw any realistic conclusions about the feasibility of this design. Wilson et al. (2017) (Wilson et al., 2017) presented Programmable VHDL Neuron Array (PVNA) - an NMA with neurons following the modified version of Message-Based Event Driven (MBED)-based neuron model(Claverol, 2000). The architecture relied on a common bus connecting PEs for neurons and synapses with time-multiplexed access. The system supported pausing the simulation for adjusting the system on-the-fly and performing external online learning with STDP. The synapse PEs were daisy-chained with modifiable connections. The implemented system supported a maximum of 100 neurons, 16 pattern generators and 200 synapses on a Xilinx Virtex-5 device and it was tested with a C. Elegans locomotor system simulation with 86 neurons, 6 pattern generators and 180 synapses. Farsa et al. (2019) (Farsa et al., 2019) provided a simplified SNN implementation with a focus on minimizing utilization figures. The authors designed it for image inference performed on 5x5 arrays of binary pixels. The neuron model was LIF with simplified operation (1.25 ms timestep and fully multiplierless). The topology was an FF-FC network (6 neurons). The implementation of a single neuron with this architecture could be clocked at 412.371 MHz.
@@
Modern implementations. Gupta et al. (2020) (Gupta et al., 2020) presented an architecture for realizing a WTA network for classification tasks. The system comprised 16 PEs, one for every neuron in the output layer of the WTA network, as well as 784 input units for the input spikes related to the pixels. It supported online learning in the form of STDP and fixed-point numerical representation of 24b width. The authors tested the system on Xilinx Virtex-6, where it was clocked at 100 MHz and achieved the inference speed of 2000 frames per second for the MNIST benchmark. Prashanth et al. (2021) (Prashanth and Ahmed, 2021) presented a general view of the implemented architecture, which we deduced to be an accelerator for a particular SCNN with LIF neurons. The architecture used an ARM core for system control and an array of PEs. The paper did not provide any concrete numbers, but we deduced that four PEs were implemented in hardware. The authors also claimed that the system supported STDP. Wang et al. (2023) (Wang et al., 2023c) presented an architecture with a specific memory optimization technique, targeting the FF SNNs implementations. They used ping-pong buffers to speed up operations. The authors focused on optimizing the sparse spike characteristics: the neuron fan-in was cut into smaller chunks of 64, which were OR’ed to see if there were any spikes incoming. The architecture supported LIF neurons and 256 PEs with one-hot encoding, which the authors claimed increased performance for high levels of connection sparsity (8.84x better performance with 90% sparsity).
3.9. Non-NMA structures related to SNN research
@@
There is a number of structures that we were not able to fairly compare with the implementations listed in Sections from 3.1 to 3.8, due to the lack of distinctive features of NMAs, as we defined them in Section 2.3. There are many such implementations reported in the literature, and we are listing two of them as examples. Palumbo et al. (2017) (Palumbo et al., 2017) presented an architecture for simulating Swarm Intelligence (SI) using SNNs,it employed a soft-core processor and a designated coprocessor to provide a systolic array for realizing partial calculations for neuronal dynamics. Shahsavari et al. (2021) presented Partially Ordered Event-Triggered System (POETS) architecture - a general-purpose parallel architecture of a number of RISC-V cores arranged in a NoC on multiple FPGAs.
@@
In many papers, the authors were focusing on optimizing parts of NMAs. As an interested reader may find information on optimizing specific their NMA designs extremely useful, we list such papers, grouped by the sub-units being optimized: spike encoding - (Gerlinghoff et al., 2022), exponential function unit - (Kim et al., 2021), STDP-based learning - (Gomar and Ahmadi, 2018; Nouri et al., 2018; Pedroni et al., 2016; Belhadj et al., 2008), ion channel dynamics unit - (Jokar and Soleimani, 2017; Mak et al., 2005), soma unit - (Pourhaj and Teng, 2010), alternative to SNN-based processing (P systems) - (Peña et al., 2019).
@@
While surveying the neuromorphic architectures, we realized that many researchers were focusing on implementing single neurons on FPGAs with hardware optimizations, accuracy of simulation and biological plausibility of the results in mind. Examples of such works, grouped by the neuron model the authors aimed at implementing, are as follows: HH - (Levi et al., 2018; Bonabi et al., 2014), LIF - (Grassia et al., 2016), AdEx - (Heidarpour et al., 2016; Gomar and Ahmadi, 2014), HR - (Heidarpur et al., 2017; Hayati et al., 2016; Kazemi et al., 2014), ML - (Hayati et al., 2015), IZH - (Pu et al., 2021; Haghiri et al., 2018), WIL - (Imani et al., 2018), Pinsky-Rinzel (PR) - (Rahimian et al., 2018), multiple - (Soleimani and Drakakis, 2018), other - (Zhang et al., 2022).
4. Trends, Tendencies, and Discussion
(a) Class population structure through years.
(b) Share of different neuron models through years.
(c) Complexity of the implemented architectures.
(d) Share of use cases for the surveyed architectures.
Figure 11. Diagrams with various properties of the classified NMAs.
[…]
Diagrams with the observed trends.
@@ 4. Trends, Tendencies, and Discussion
We surveyed 129 papers describing NMAs that provided enough information for us to apply the Taxonomy rules and derive further metrics. Additional 27 papers consisted of non-NMA structures and NMA-related research, as described in Section 3.9. Around 86.78% of architecture featured a parallel architecture, with 30.58% being fully parallel. Some form of online learning was featured in 31.4% of implementations, which is a rather small number - most likely due to the lack of agreement on how to realize brain-like learning in hardware efficiently. As Figure 11(a) suggests, the NMAs on FPGAs have been gaining popularity throughout the years, with the most implementations presented around the year 2021. We can see that the three most populous classes are Class 2 (38), Class 4 (22) and Class 6 (17). This is unsurprising as the collocated memory and computation is the easiest Trait to include, and thus we can see them appearing often. Class 7 consists of only 10 examples, but they can also be seen throughout the years, showing interest in implementing fully neuromorphic systems. Close to 53.5% of the NMAs realized up to and including 1000 neurons. This can be due to the classification tasks being the most popular use case – 40% of NMAs according to the Figure 11(d) – and those do not require large networks. The target benchmarks should be thus reconsidered and new ones should be introduced to test more complex NMAs. Moreover, 28% of architectures target neuroscience, reinforcing the close connection between biological and spiking neural networks. The most popular neuron model is LIF (47.29%), followed by IZH (18.6%), IF (14.7%) and HH (9.3%) - the rest of the models were utilized in around 10% of the architectures. As LIF neurons are the simplest to implement, IZH give similar spiking behaviors with reduced complexity to the complete and HH neurons are popular for neuroscientific experiments - those results are not surprising.
@@
Figure 12. Predicted LB number required to implement $10^{11}$ of neurons on an FPGA by Class.
@@
The hardware constraints were a crucial factor when selecting the appropriate neuron model and not many researchers decided to implement the complex HH model. The most popular topology is FF-FC (33.3%), then SCNN (12.4%), A2A (10.9%), biology-related ”fixed” topologies (8.5%), WTA (7%) and LSMs (2.3%) - other topologies were usually special cases with a single representative. Once again, the influence of the use case in topology choice is rather apparent. Interestingly, a majority of NMAs (86%) were implemented on Xilinx/AMD devices, which may suggest that FPGA-based NMAs development is dependent on a single manufacturer when it comes for creating larger and faster implementations. Moreover, nearly 77% of the architectures utilized at most 50% of the available resources, regardless of the platform size and manufacturer, and there were only five examples of systems clocked at above 200 MHz, which suggests that even today the FPGAs are not utilized to their full potential in the field of NMA development. Figure 12 shows the trend of available logic resources on FPGAs, based on the size of the AMD/Xilinx Virtex FPGA device family [^11] The newest devices from AMD - Versal - are discarded as they are not pure FPGAs, but rather System-On-a-Chip devices with relatively small FPGA fabric (as a percentage of the entire chip). (black dots) through the years and required logic resources for implementing $10^{11}$ neurons – the estimated size of the human brain – on a single chip for the most promising architectures from every Class. We can see that, if we discard the impact of embedded memory size and improvements in on-chip communication and assume this trend to hold, Classes 0 and 3 can be expected to achieve human-brain scale around the year 2035 as they require the least amount of on-chip logic resources and scale the best in single-chip implementations. The rest need more time – Classes 1, 2, 4, 6 and 7 would achieve the goal after 2055 and Class 5 (prediction is rather vague here, because it is based on a system that used multiple FPGAs – BiCoSS) would achieve it after 2045. They would benefit more from adding more chips, as it is already being done(Kudithipudi et al., 2025). However, considering that IEEE suggests in their International Roadmap for Devices and Systems (IRDS) from 2022 that Moore’s Law will be upheld for additional 10-15 years(IEEE International Roadmap for Devices and Systems, 2022), only Classes 0 and 3 seem likely to achieve the aforementioned goal in the predicted years.
5. Conclusion

### v3

@@ 3.8 Class 7 (Traits: fully parallel, collocated computation and memory, asynchronous network update)
Early implementations. Allen et al. (2005) [189] presented a platform for pattern recognition in olfactory applications. The system supported IF neuron model with weights uniformly set to 1, so the output spike train of the neuron had a firing rate equal to the average rate of the input spike trains. We deduced that architecture-wise the system was an array of PEs - one for every neuron. The authors reported eight neurons with 1024 input synapses implemented on the Xilinx Virtex-II Pro device. Girau et al. (2006) [190] presented an architecture for a Local Excitatory Global Inhibitory Oscillator Network (LEGION)[191], based on IF neurons, for image inference.
Figure 10: Simplified diagram of a Class 7 NMA.
@@
Architecture-wise, due to the inherent nature of the LEGION-type SNNs, the implementation comprised a 2D array of PEs. Neurons were interconnected with excitatory connections, and there was a single global inhibitor connected to all neurons. The size of the array was equal in size to the size of the image under classification. Cassidy et al. (2007) [192] presented a 2D array of LIF neurons capable of realizing Spatio-Temporal Receptive Fields (STRFs), neural parameter optimization algorithms and STDP. The architecture comprised 32 PEs set up as an array, connected to the weight and delay memory. The datapath was time-multiplexed to allow for parallel processing and serial communication. Every neuron was connected to its four neighbors. The system was implemented on Xilinx Spartan-3 device and was clocked at 50 MHz.
@@
Early modern implementations. Bonabi et al. (2021) [193] the authors presented a network of 16 HH neurons on a Xilinx Virtex-7 device. The neuronal dynamics were approximated with the CORDIC algorithm and step-by-step time integration. The authors focused on neuronal pools, i.e., groups of interconnected neurons (in A2A fashion) with specific functions, which could be either diffusive or localized. Dean et al. (2014) [194] presented DANNA - an NMA that comprised an array of multi-purpose elements that could be configured to perform a function of a neuron, synapse or fan-in/fan-out connection. The architecture was a 2D grid of elements, which be neurons, synapses or fan-out elements. The architecture supported the LIF neuron model with axonal delays. The synapse model supported long-term potentiation and depression for the weights. The network topology was obtained by performing the genetic algorithm for finding the optimal implementation. The authors tested the system [^10] There were many examples of the use cases for this architecture, including NeoN[195], which was a neuromorphic control for autonomous robotic navigation. with a randomly connected network of 10000 neurons, implemented on a Xilinx Virtex-7 device with a 9b fixed-point number representation, clocked at eight megahertz, while the elements was clocked at 64 MHz. The system was also updated in 2018 and presented as DANNA 2[196]. Nanami et al. (2016) [197] presented a modification of an existing Digital Spiking Silicon Neuron (DSSN)[198] idea. The proposed architecture used 16 DSSNs connected to 16 synapse units, each supporting 256 external synapses - all operating at 10x real-time speed. The number representation was 18b. There were no implementation details provided in the paper, so it is difficult to draw any realistic conclusions about the feasibility of this design. Wilson et al. (2017) [199] presented Programmable VHDL Neuron Array (PVNA) - an NMA with neurons following the modified version of Message-Based Event Driven (MBED)-based neuron model[56]. The architecture relied on a common bus connecting PEs for neurons and synapses with time-multiplexed access. The system supported pausing the simulation for adjusting the system on-the-fly and performing external online learning with STDP. The synapse PEs were daisy-chained with modifiable connections. The implemented system supported a maximum of 100 neurons, 16 pattern generators and 200 synapses on a Xilinx Virtex-5 device and it was tested with a C. Elegans locomotor system simulation with 86 neurons, 6 pattern generators and 180 synapses. Farsa et al. (2019) [200] provided a simplified SNN implementation with a focus on minimizing utilization figures. The authors designed it for image inference performed on 5x5 arrays of binary pixels. The neuron model was LIF with simplified operation (1.25 ms timestep and fully multiplierless). The topology was an FF-FC network (6 neurons). The implementation of a single neuron with this architecture could be clocked at 412.371 MHz.
@@
Modern implementations. Gupta et al. (2020) [201] presented an architecture for realizing a WTA network for classification tasks. The system comprised 16 PEs, one for every neuron in the output layer of the WTA network, as well as 784 input units for the input spikes related to the pixels. It supported online learning in the form of STDP and fixed-point numerical representation of 24b width. The authors tested the system on Xilinx Virtex-6, where it was clocked at 100 MHz and achieved the inference speed of 2000 frames per second for the MNIST benchmark. Prashanth et al. (2021) [202] presented a general view of the implemented architecture, which we deduced to be an accelerator for a particular SCNN with LIF neurons. The architecture used an ARM core for system control and an array of PEs. The paper did not provide any concrete numbers, but we deduced that four PEs were implemented in hardware. The authors also claimed that the system supported STDP. Wang et al. (2023) [203] presented an architecture with a specific memory optimization technique, targeting the FF SNNs implementations. They used ping-pong buffers to speed up operations. The authors focused on optimizing the sparse spike characteristics: the neuron fan-in was cut into smaller chunks of 64, which were OR’ed to see if there were any spikes incoming. The architecture supported LIF neurons and 256 PEs with one-hot encoding, which the authors claimed increased performance for high levels of connection sparsity (8.84x better performance with 90% sparsity).
3.9 Non-NMA structures related to SNN research
@@
There is a number of structures that we were not able to fairly compare with the implementations listed in Sections from 3.1 to 3.8, due to the lack of distinctive features of NMAs, as we defined them in Section 2.3. There are many such implementations reported in the literature, and we are listing two of them as examples. Palumbo et al. (2017) [204] presented an architecture for simulating Swarm Intelligence (SI) using SNNs,it employed a soft-core processor and a designated coprocessor to provide a systolic array for realizing partial calculations for neuronal dynamics. Shahsavari et al. (2021) presented Partially Ordered Event-Triggered System (POETS) architecture - a general-purpose parallel architecture of a number of RISC-V cores arranged in a NoC on multiple FPGAs.
@@
In many papers, the authors were focusing on optimizing parts of NMAs. As an interested reader may find information on optimizing specific their NMA designs extremely useful, we list such papers, grouped by the sub-units being optimized: spike encoding - [205], exponential function unit - [206], STDP-based learning - [207, 208, 209, 210], ion channel dynamics unit - [211, 212], soma unit - [213], alternative to SNN-based processing (P systems) - [214].
@@
While surveying the neuromorphic architectures, we realized that many researchers were focusing on implementing single neurons on FPGAs with hardware optimizations, accuracy of simulation and biological plausibility of the results in mind. Examples of such works, grouped by the neuron model the authors aimed at implementing, are as follows: HH - [215, 216], LIF - [217], AdEx - [218, 219], HR - [220, 221, 222], ML - [223], IZH - [224, 225], WIL - [226], Pinsky-Rinzel (PR) - [227], multiple - [228], other - [229].
4 Trends, Tendencies, and Discussion
@@
We surveyed 129 papers describing NMAs that provided enough information for us to apply the Taxonomy rules and derive further metrics. Additional 27 papers consisted of non-NMA structures and NMA-related research, as described in Section 3.9. Around 86.78% of architecture featured a parallel architecture, with 30.58% being fully parallel. Some form of online learning was featured in 31.4% of implementations, which is a rather small number - most likely due to the lack of agreement on how to realize brain-like learning in hardware efficiently. As Figure 11 suggests, the NMAs on FPGAs have been gaining popularity throughout the years, with the most implementations presented around the year 2021. We can see that the three most populous classes are Class 2 (38), Class 4 (22) and Class 6 (17). This is unsurprising as the collocated memory and computation is the easiest Trait to include, and thus we can see them appearing often. Class 7 consists of only 10 examples, but they can also be seen throughout the years, showing interest in implementing fully neuromorphic systems. Close to 53.5% of the NMAs realized up to and including 1000 neurons. This can be due to the classification tasks being the most popular use case – 40% of NMAs according to the Figure 14 – and those do not require large networks. The target benchmarks should be thus reconsidered and new ones should be introduced to test more complex NMAs. Moreover, 28% of architectures target neuroscience, reinforcing the close connection between biological and spiking neural networks. The most popular neuron model is LIF (47.29%), followed by IZH (18.6%), IF (14.7%) and HH (9.3%) - the rest of the models were utilized in around 10% of the architectures. As LIF neurons are the simplest to implement, IZH give similar spiking behaviors with reduced complexity to the complete and HH neurons are popular for neuroscientific experiments - those results are not surprising. The hardware constraints were a crucial factor when selecting the appropriate neuron model and not many researchers decided to implement the complex HH model. The most popular topology is FF-FC (33.3%), then SCNN (12.4%), A2A (10.9%), biology-related ”fixed” topologies (8.5%), WTA (7%) and LSMs (2.3%) - other topologies were usually special cases with a single representative. Once again, the influence of the use case in topology choice is rather apparent. Interestingly, a majority of NMAs (86%) were implemented on Xilinx/AMD devices, which may suggest that FPGA-based NMAs development is dependent on a single manufacturer when it comes for creating larger and faster implementations. Moreover, nearly 77% of the architectures utilized at most 50% of the available resources, regardless of the platform size and manufacturer, and there were only five examples of systems clocked at above 200 MHz, which suggests that even today the FPGAs are not utilized to their full potential in the field of NMA development.
Figure 11: Class population structure through years.
Figure 12: Share of different neuron models through years.
Figure 13: Complexity of the implemented architectures.
Figure 14: Share of use cases for the surveyed architectures.
@@
Figure 15: Predicted LB number required to implement $10^{11}$ of neurons on an FPGA by Class.
@@
Figure 15 shows the trend of available logic resources on FPGAs, based on the size of the AMD/Xilinx Virtex FPGA device family [^11] The newest devices from AMD - Versal - are discarded as they are not pure FPGAs, but rather System-On-a-Chip devices with relatively small FPGA fabric (as a percentage of the entire chip). (black dots) through the years and required logic resources for implementing $10^{11}$ neurons – the estimated size of the human brain – on a single chip for the most promising architectures from every Class. We can see that, if we discard the impact of embedded memory size and improvements in on-chip communication and assume this trend to hold, Classes 0 and 3 can be expected to achieve human-brain scale around the year 2035 as they require the least amount of on-chip logic resources and scale the best in single-chip implementations. The rest need more time – Classes 1, 2, 4, 6 and 7 would achieve the goal after 2055 and Class 5 (prediction is rather vague here, because it is based on a system that used multiple FPGAs – BiCoSS) would achieve it after 2045. They would benefit more from adding more chips, as it is already being done[230]. However, considering that IEEE suggests in their International Roadmap for Devices and Systems (IRDS) from 2022 that Moore’s Law will be upheld for additional 10-15 years[231], only Classes 0 and 3 seem likely to achieve the aforementioned goal in the predicted years.
5 Conclusion